# Synthetic Tabular Data Validation: A Divergence-Based Approach

### v1

@@ 4 Experiments
We aim to validate the effectiveness of our proposed divergence estimator for data generation. We achieve this through an exhaustive experimental evaluation in four different experiments, described in Table 1 designed to assess the estimator’s performance under various conditions. The initial set of experiments focuses on controlled scenarios that involve comparisons of simple theoretical data distributions. Subsequently, the complexity is gradually increased by incorporating generative scenarios that reproduce real-world applications.
@@
We comprehensively analyze the influence of sample size on the estimator’s performance. This analysis covers the impact of the size of the training and validation set for the discriminator network and the number of samples used during the generative process. To gain a deeper understanding, we evaluated different configurations for each parameter, as the table shows. For every experiment, all possible combinations of $N$, $M$, and $L$ are executed 5 times with unique random seeds. This rigorous approach employing multiple random seeds mitigates the effects of inherent randomness within training processes. Consequently, it provides a statistically robust evaluation, providing a more reliable representation of the observed trends. Ultimately, this comprehensive analysis facilitates identifying the configuration that yields the most accurate assessment of the divergence between data distributions.
@@
For complete transparency and reproducibility, the data and code used in this study are publicly available on our repository.
@@ 4.1 Experiment 1
Table 4.1 presents the results obtained for $D_{\mathbb{KL}}$ and $D_{\mathbb{JS}}$ between random multivariate Gaussian distributions. A significant advantage of this study is the availability of a closed-form solution for the KL divergence, which serves as the ground truth value. For the JS divergence, we use the MC simulation estimate as the ground truth. To ensure the reliability of this experiment, we opted for multivariate Gaussian distributions with a dimensionality of 10. This allows us to compute the true divergence values and assess the accuracy of estimating the ratio of the proposed discriminator. The results reveal a critical relationship between sample size and estimation accuracy, particularly for the discriminator-based approach. When the number of samples is limited ($M=20$, $L=20,200$), the estimated divergence from the discriminator deviates significantly from both the analytical value and the MC estimate. This suggests a potential overfitting due to insufficient data, leading to an underestimation of the true divergence. However, as the number of samples increases, the estimated divergence from the discriminator network progressively converges towards the ground truth value, followed by improved confidence interval precision. Figs. 3 and 4 confirm this trend. Both figures illustrate how increasing the number of training samples $M$ and validation samples $L$ reduces the error associated with the estimated divergence ratio. Further support for the correlation between sample size and estimation accuracy comes from the discriminator loss curves depicted in Fig. 5. As detailed in [21], the loss function converges to a constant value approximating the Jensen-Shannon divergence. The figure visually confirms this concept, as the loss curves flatten with increasing training $M$ and validation $L$ samples. In contrast, a low number of samples (e.g. $M=20$, $L=20$) results in larger fluctuations in the loss function, potentially indicating discriminator overfitting. These validations demonstrate the effectiveness of our proposed method: with sufficient training data, the discriminator can accurately learn the density ratio and provide reliable estimates of the divergences, particularly for the JS divergence.
@@
Figure 3: Estimation error representation for $D_{\mathbb{KL}}$. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
Figure 4: Estimation error representation for $D_{\mathbb{JS}}$. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
@@
Leveraging Experiment 1, we further emphasize the superior robustness of our proposed approach in estimating the JS divergence, regardless of variations in distribution separation. Fig. 6 illustrates the relation between estimated divergences and ground truth values as the separation between two 4-dimensional multivariate Gaussian distributions increases. As observed, the relation for the JS divergence remains almost linear until it approaches 1. While the relationship appears less linear for values near 1, it is important to note that the primary application of this discriminator lies in comparing similar distributions. Therefore, the focus should be placed on the performance for low JS divergences. Meanwhile, the KL divergence exhibits a nonlinear relation starting from the first comparisons, with the estimated KL deviating from the ground truth probably because of bias introduced by the discriminator.
Figure 6: Relation between ground truth and estimated divergences as distribution separation increases.
@@ 4.2 Experiment 2
This experiment investigates the performance of the proposed divergence estimation method in Gaussian mixture distributions. These distributions offer a higher level of complexity compared to those used in Experiment 1. Each Gaussian mixture distribution comprises two independent isotropic Gaussian components with distinct mixing probabilities (weights assigned to each component within the mixture). Unlike the prior case, where analytical ground truth for divergence was obtainable, here and in the following experiment, we rely solely on MC approximations as ground truth. Similar results were obtained for this experiment compared to the first. Tables 4 and 5 focus on a practical range for training sample sizes $M$ and validation sample sizes $L$ (both set to $200,2000$) and present the estimated divergences obtained using our proposed method compared to MC estimates. The results demonstrate that, for both divergences, the proposed method achieves results comparable to the MC estimates. These findings further support the previously observed correlation between sample size and estimation accuracy. Additionally, for both divergence measures, we observe a narrowing of confidence intervals with increasing sample sizes. This trend indicates that the estimation becomes progressively more precise as the amount of available data increases. Overall, the results from this experiment underscore the effectiveness of the proposed method in estimating divergences for more intricate distributions, such as Gaussian mixtures. The remaining analysis can be found in the Appendix.
@@
Figure 7: Estimated $D_{\mathbb{KL}}$ and $D_{\mathbb{JS}}$ compared to ground truth values for varying sample sizes of $N$. Note the independent y-axes for each divergence measure due to their inherent scale differences, with $D_{\mathbb{JS}}$ exhibiting lower values than $D_{\mathbb{KL}}$.
@@ 4.3 Experiment 3
The following experiments introduce the concept of a generative process, where synthetic data are generated based on $N$ samples from a certain data distribution. We investigate the influence of the quality of the GM in terms of the number of samples used to generate. This specific experiment uses a Gaussian mixture with two components as the real distribution. We employ a Gaussian mixture model (GMM) as the GM. The trained GMM then generates synthetic data that serve as an approximation of the real data. Finally, we estimate the divergence using our approach with a sufficient number of samples to achieve a low confidence interval and compare it to the MC simulation estimation. We analyze the behavior of the GMM under different training configurations, varying the number of training samples $N$. Since previous experiments demonstrated that a high number of samples for both $M$ and $L$ achieve better divergence estimations, for this experiment we have fixed $M=2000$ and $L=2000$. We compare the estimated divergence errors for this combination of training samples $M$ and validation $L$ used by the divergence estimator. Fig. 7 shows the estimated divergences as $N$ increases. The results demonstrate a clear correlation between the size of the training data for the GMM, $N$, and the estimated divergence errors. When the GMM is poorly trained due to a limited number of samples ($N=10$), it is unable to effectively capture the underlying data patterns. This leads to generated data that significantly deviate from the real distribution. Consequently, the real divergence values increase significantly and a large estimation error occurs, particularly for unbounded $D_{\mathbb{KL}}$. $D_{\mathbb{JS}}$, which is inherently bounded between 0 and 1, exhibits less extreme error values. As the number of training samples for the GMM increases ($N=200$ and $N=2000$), the generated data become more representative of the real distribution. This results in lower divergence values and reduced estimation errors for $D_{\mathbb{KL}}$ and $D_{\mathbb{JS}}$. This finding supports the validity of our proposed validation technique for generative processes involving Gaussian mixture distributions.
@@ 4.4 Experiment 4
The final experiment introduces a more realistic setting by incorporating a real-world dataset along with its synthetic counterpart generated using a GM. This scenario retains the concept of a generative process, but the target distribution becomes the real-world data itself. The dataset chosen is Adult. As described in [24], it contains information on $32,561$ individuals used to predict their annual income exceeding \$50,000. It represents an extract from the $1994$ U.S. Census database [25]. For our experiments and after the previous analysis done, we used a subset of $10,000$ samples obtained from [26]. This dataset comprises 14 features encompassing various data types, including categorical, binary, and integer values. Based on the previous analysis of the influence of the sample size, we fixed the hyperparameters for this experiment at $N=10000$, $M=7500$, and $L=1000$. To evaluate GMs, we compared the performance of two state-of-the-art approaches: the widely used CTGAN[12] and the VAE-based generator proposed in [23] (VAE), which achieved superior results based on several validation criteria. Table 6 confirms this, obtaining lower divergence results for [23]. This experiment focuses on evaluating the ability of the divergence estimator to assess the similarity between the real data and its synthetic counterpart. This experiment provides valuable insights into the practical application of the proposed method in real-world scenarios where measuring the similarity between real and synthetic data is crucial.
@@ 5 Conclusion
This research proposes a novel and effective approach for validating synthetic tabular data generated by various models. The core of this method lies in using a divergence estimator based on a probabilistic classifier to capture the discrepancies between the real and synthetic data distributions. This approach overcomes the limitations associated with traditional marginal divergence comparisons by considering the joint distribution. While marginal comparisons assess the similarity of individual features between real and synthetic data, they can be misleading. Even if the marginal distributions of each feature appear similar, the joint distribution, which captures the relationships between features, may differ significantly. This can lead to unrealistic synthetic data, where individual features appear plausible but their co-occurrences are not representative of the real data. By considering the joint distribution, our proposed method provides a more comprehensive assessment of the quality of synthetic tabular data.
@@
This research offers significant contributions that extend beyond the specific tabular data validation domain. The proposed methodology facilitates better validation practices for various fields dependent on GMs. Its key strength lies in the capture of complex relationships between different data distributions, leading to more robust and reliable validation processes. In addition to the positive results, the study also highlights the importance of the quality of the GM. When it is inadequately trained due to insufficient data, it can significantly impact the accuracy of the divergence estimation. This emphasizes the need to carefully consider GM training procedures to ensure reliable validation results.
Several promising avenues exist for future research. One direction involves exploring the potential to extend the proposed approach to more complex data structures, such as images or time series data. Additionally, investigating the integration of this validation technique within GM training pipelines could enable the development of self-improving GMs that can automatically adjust their parameters to generate data that closely resemble the target distribution. Finally, developing a robust methodology to estimate divergences when limited samples are available presents a crucial challenge. Addressing this challenge would further enhance the versatility and practicality of the proposed validation technique in real-world scenarios with restricted data availability. Furthermore, it would be beneficial to investigate alternative density ratio estimation techniques, such as those presented in [27, 28]. These methods may offer advantages in terms of sample size requirements, estimation error, or other relevant properties.
@@ Acknowledgments
This research was supported by GenoMed4All and SYNTHEMA projects. Both have received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 101017549 and 101095530, respectively. The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@
Figure 8: Estimation error representation for $D_{\mathbb{KL}}$. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
Figure 9: Estimation error representation for $D_{\mathbb{JS}}$. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
Figure 10: Discriminator loss curves for Use Case 1. The loss curves show a clear overfitting due to low sample sizes. Green and red dashed lines represent theoretical convergence values.
Building on Experiment 1, Experiment 2 uses the same validation process. Table 7 expands on previous findings by presenting the performance of our approach with a limited number of training and validation samples for the divergence estimator. The results demonstrate convergence towards the ground truth divergence values as the number of samples used for both training and validation increases. This trend is further corroborated by Figs. 8 and 9. Additionally, Fig. 10 explores the behavior of the discriminator’s loss function for various combinations of sample sizes.

### v2

@@ 4 Experiments
We aim to validate the effectiveness of our proposed divergence estimator for data generation. We achieve this through an exhaustive experimental evaluation in four different experiments, described in Table 1 designed to assess the estimator’s performance under various conditions. The initial set of experiments focuses on controlled scenarios that involve comparisons of simple theoretical data distributions. Subsequently, the complexity is gradually increased by incorporating generative scenarios that reproduce real-world applications. This progression in complexity not only tests the estimator’s ability to handle increasingly challenging conditions but also demonstrates its scalability and reliability. The increasing complexity is essential to ensure that our estimator remains accurate and efficient, even as the distributional assumptions become less defined and more reflective of the irregularities typically found in real data.
@@
We comprehensively analyze the influence of sample size on the estimator’s performance. This analysis covers the impact of the training and validation sets sizes for the discriminator network and the number of samples used during the generative process. To gain a deeper understanding, we evaluated different configurations for each parameter, as the table shows. For every experiment, all possible combinations of $N$, $M$, and $L$ are executed five times with unique random seeds. This rigorous approach employing multiple random seeds mitigates the effects of inherent randomness within training processes. Consequently, it provides a statistically robust evaluation and a more reliable representation of the observed trends. Ultimately, this comprehensive analysis facilitates identifying the configuration that yields the most accurate assessment of the divergence between data distributions.
@@
In the experiments conducted, the parameter values of the different models used were explicitly chosen to yield illustrative results to analyze the performance of the proposed estimator in validating synthetic data generation. We recognize that alternative parameters could be employed or fine-tuning could be undertaken to identify optimal settings; however, we intend to observe the effect of the input data on the models as such and to demonstrate their potential. This approach allows us to focus on understanding how variations in input data influence the estimator’s ability to validate synthetic datasets, thereby showcasing the practical applicability and robustness of the estimator across different data scenarios.
For complete transparency and reproducibility, the data and code used in this study are publicly available on our repository (https://github.com/Patricia-A-Apellaniz/divergence_estimator).
@@ 4.1 Experiment 1
Table 4.1 presents the results obtained for $D_{\mathbb{KL}}$ and $D_{\mathbb{JS}}$ between random multivariate Gaussian distributions. A significant advantage of this study is the availability of a closed-form solution for the KL divergence, which serves as the ground truth value. We use the MC simulation estimate for the JS divergence as the ground truth. To ensure the reliability of this experiment, we opted for multivariate Gaussian distributions with a dimensionality of 10. This allows us to compute the true divergence values and assess the accuracy of estimating the ratio of the proposed discriminator. The results reveal a critical relation between sample size and estimation accuracy, particularly for the discriminator-based approach. When the number of samples is limited ($M=20$, $L=20,200$), the estimated divergence from the discriminator deviates significantly from both the analytical value and the MC estimate. This suggests a potential overfitting due to insufficient data, leading to underestimating the true divergence. However, as the number of samples increases, the estimated divergence from the discriminator network progressively converges towards the ground truth value, followed by improved confidence interval precision. Figs. 3 and 4 confirm this trend. Both figures illustrate how increasing the number of training samples $M$ and validation samples $L$ reduces the error associated with the estimated divergence ratio. Further support for the correlation between sample size and estimation accuracy comes from the discriminator loss curves depicted in Fig. 5. As detailed in [22], the loss function converges to a constant value approximating the Jensen-Shannon divergence. The figure visually confirms this concept, as the loss curves flatten with increasing training $M$ and validation $L$ samples. In contrast, a low number of samples (e.g., $M=20$, $L=20$) results in larger fluctuations in the loss function, potentially indicating discriminator overfitting. These validations demonstrate the effectiveness of our proposed method: with sufficient training data, the discriminator can accurately learn the density ratio and provide reliable estimates of the divergences, particularly for the JS divergence.
@@
Figure 3: Estimation error representation for $D_{\mathbb{KL}}$ in Experiment 1. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
Figure 4: Estimation error representation for $D_{\mathbb{JS}}$ in Experiment 1. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
@@
Figure 6: Relation between ground truth and estimated divergences as distribution separation increases in Experiment 1.
Leveraging Experiment 1, we further emphasize the superior robustness of our proposed approach in estimating the JS divergence, regardless of variations in distribution separation. Fig. 6 illustrates the relation between estimated divergences and ground truth values as the separation between two 4-dimensional multivariate Gaussian distributions increases. As observed, the relation for the JS divergence remains almost linear until it approaches 1. While the relation appears less linear for values near 1, it is essential to note that the primary application of this discriminator lies in comparing similar distributions. Therefore, the focus should be on the performance of low JS divergences. Meanwhile, the KL divergence exhibits a nonlinear relation starting from the first comparisons, with the estimated KL deviating from the ground truth probably because of bias introduced by the discriminator.
@@ 4.2 Experiment 2
This experiment investigates the performance of the proposed divergence estimation method in Gaussian mixture distributions. These distributions offer a higher complexity level than those used in Experiment 1. Each Gaussian mixture distribution comprises two independent isotropic Gaussian components with distinct mixing probabilities (weights assigned to each component within the mixture). Unlike the prior case, where analytical ground truth for divergence was obtainable, we rely solely on MC approximations as ground truth here and in the following experiment. Similar results were obtained for this experiment compared to the first.
Tables 4 and 5 focus on a practical range for training sample sizes $M$ and validation sample sizes $L$ (both set to $200,2000$) and present the estimated divergences obtained using our proposed method compared to MC estimates. The results demonstrate that, for both divergences, the proposed method achieves results comparable to the MC estimates. These findings further support the previously observed correlation between sample size and estimation accuracy. Additionally, we observe a narrowing of confidence intervals with increasing sample sizes for both divergence measures. This trend indicates that the estimation becomes progressively more precise as the amount of available data increases. Overall, this experiment’s results underscore the proposed method’s effectiveness in estimating divergences for more intricate distributions, such as Gaussian mixtures. The remaining analysis can be found in the Appendix.
@@ 4.3 Experiment 3
Figure 7: Estimated $D_{\mathbb{KL}}$ and $D_{\mathbb{JS}}$ compared to ground truth values for varying sample sizes of $N$ in Experiment 3. Note the independent y-axes for each divergence measure due to their inherent scale differences, with $D_{\mathbb{JS}}$ exhibiting lower values than $D_{\mathbb{KL}}$.
The following experiments introduce the concept of a generative process, where synthetic data are generated based on $N$ samples from a particular data distribution. We investigate the influence of the quality of the GM in terms of the number of samples used to generate. This specific experiment uses a Gaussian mixture with two components as the real distribution. We employ a Gaussian mixture model (GMM) as the GM. The trained GMM then generates synthetic data that serve as an approximation of the real data. Finally, we estimate the divergence using our approach with a sufficient number of samples to achieve a low confidence interval and compare it to the MC simulation estimation. We analyze the behavior of the GMM under different training configurations, varying the number of training samples $N$. Since previous experiments demonstrated that a high number of samples for both $M$ and $L$ achieve better divergence estimations, we have fixed $M=2000$ and $L=2000$ for this experiment. We compare the estimated divergence errors for this combination of training samples $M$ and validation $L$ used by the divergence estimator.
Fig. 7 shows the estimated divergences as $N$ increases. The results demonstrate a clear correlation between the size of the training data for the GMM, $N$, and the estimated divergence errors. When the GMM is poorly trained due to a limited number of samples ($N=10$), it cannot effectively capture the underlying data patterns. This leads to generated data that significantly deviate from the real distribution. Consequently, the real divergence values increase significantly, and a large estimation error occurs, particularly for unbounded $D_{\mathbb{KL}}$. $D_{\mathbb{JS}}$, which is inherently bounded between 0 and 1, exhibits less extreme error values. As the number of training samples for the GMM increases ($N=200$ and $N=2000$), the generated data become more representative of the real distribution. This results in lower divergence values and reduced estimation errors for $D_{\mathbb{KL}}$ and $D_{\mathbb{JS}}$. This finding supports the validity of our proposed validation technique for generative processes involving Gaussian mixture distributions.
@@ 4.4 Experiment 4
The final experiment introduces a more realistic setting by incorporating a real-world dataset along with its synthetic counterpart generated using a GM. This scenario retains the concept of a generative process, but the target distribution becomes the real-world data itself. The dataset chosen is Adult. As described in [26], it contains information on $32,561$ individuals used to predict their annual income exceeding \$50,000. It represents an extract from the $1994$ U.S. Census database [27]. For our experiments and after the previous analysis was done, we used a subset of $10,000$ samples obtained from [28]. This dataset comprises 14 features encompassing various data types, including categorical, binary, and integer values. Based on the previous analysis of the influence of the sample size, we fixed the hyperparameters for this experiment at $N=10000$, $M=7500$, and $L=1000$. To evaluate GMs, we compared the performance of two state-of-the-art approaches: the widely used CTGAN[12] and the VAE-based generator proposed in [25] (VAE), which achieved superior results based on several validation criteria.
Table 6 confirms this, obtaining lower divergence results for [25]. To explain why the VAE may outperform the CTGAN in our experiments, we refer to [25], which compares the VAE generator with CTGAN and TVAE using metrics similar to ours. The results support our observations. We suggest that VAEs perform better due to their stability in managing complex data distributions, unlike GANs, which struggle with convergence. This stability in VAEs contributes to their enhanced ability to generate more realistic synthetic data, highlighting significant differences in model architecture and distribution handling. This experiment focuses on evaluating the ability of the divergence estimator to assess the similarity between the real data and its synthetic counterpart. This experiment provides valuable insights into the practical application of the proposed method in real-world scenarios where measuring the similarity between real and synthetic data is crucial.
@@ 5 Conclusion
This research proposes a novel and practical approach for validating synthetic tabular data generated by various models. The core of this method lies in using a divergence estimator based on a probabilistic classifier to capture the discrepancies between the real and synthetic data distributions. This approach overcomes the limitations associated with traditional marginal divergence comparisons by considering the joint distribution. While marginal comparisons assess the similarity of individual features between real and synthetic data, they can be misleading. Even if the marginal distributions of each feature appear similar, the joint distribution, which captures the relations between features, may differ significantly. This can lead to unrealistic synthetic data, where individual features appear plausible, but their co-occurrences do not represent the real data. By considering the joint distribution, our proposed method provides a more comprehensive assessment of the quality of synthetic tabular data.
@@
This research offers significant contributions that extend beyond the specific tabular data validation domain. The proposed methodology facilitates better validation practices for various fields dependent on GMs. Its key strength lies in capturing complex relations between different data distributions, leading to more robust and reliable validation processes. In addition to the positive results, the study also highlights the importance of the quality of the GM. When it is inadequately trained due to insufficient data, it can significantly impact the accuracy of the divergence estimation. This emphasizes the need to consider GM training procedures carefully to ensure reliable validation results.
Several promising avenues exist for future research. One direction involves exploring the potential to extend the proposed approach to more complex data structures, such as images or time series data. Additionally, investigating the integration of this validation technique within GM training pipelines could enable the development of self-improving GMs that can automatically adjust their parameters to generate data that closely resemble the target distribution. Moreover, addressing the impact of changing the assumption of prior probabilities for the source distributions to be compared, where the number of real and synthetic samples may vary, presents a significant challenge. Techniques to mitigate the effects of unbalanced classes in the regression framework could be explored. These techniques may include adaptive sampling strategies, class weighting, or methods specifically tailored to handle imbalanced data distributions. Implementing such approaches could enhance the robustness and applicability of the validation technique across diverse datasets and real-world applications. Finally, developing a robust methodology to estimate divergences when limited samples are available presents a crucial challenge. Addressing this challenge would further enhance the versatility and practicality of the proposed validation technique in real-world scenarios with restricted data availability. Furthermore, it would be beneficial to investigate alternative density ratio estimation techniques, such as those presented in [29, 30]. These methods may offer advantages regarding sample size requirements, estimation error, or other relevant properties.
@@ Acknowledgments
This research was supported by GenoMed4All and SYNTHEMA projects. Both have received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 101017549 and 101095530, respectively. The authors declare that they have no known competing financial interests or personal relations that could have appeared to influence the work reported in this paper.
@@
Leveraging the validation process established in Experiment 1, Experiment 2 employs the same methodology. The remaining analyses for Experiment 2 are presented as follows.
Table 7 extends the findings of the previous experiment by investigating the performance of our proposed approach with a restricted number of training and validation samples for the divergence estimator. The results indicate a trend of convergence towards the known ground truth divergence values as the number of samples utilized for both training and validation is progressively increased.
Figs. 8 and 9 provide further substantiation for this observed trend.
Additionally, Fig. 10 delves into the behavior of the discriminator’s loss function across various sample size combinations.
@@
Figure 8: Estimation error representation for $D_{\mathbb{KL}}$ in Experiment 2. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
Figure 9: Estimation error representation for $D_{\mathbb{JS}}$ in Experiment 2. Results are shown for different combinations of training sample sizes $M$ and validation sample sizes $L$. As expected, a decrease and precision in the error is observed with increasing values of $M$ and $L$.
Figure 10: Discriminator loss curves for Experiment 2. The loss curves show a clear overfitting due to low sample sizes. Green and red dashed lines represent theoretical convergence values.